OpenStudy (anonymous):

Please, help (NO ANSWERS, PLEASE :) Choose the equation below that represents the line passing through the point (−2, −3) with a slope of −6. y + 3 = −6(x + 2) y + 2 = 6(x + 3) y − 3 = −6(x − 2) y − 2 = 6(x − 3)

OpenStudy (radar):

Hint: It must also have a slope of -6 or x will have a coefficient of -6 when placed in slope-intercept form.

OpenStudy (radar):

Also when the value of -2 is substituted for x, y will equal -3).

OpenStudy (anonymous):

so C?

OpenStudy (radar):

Just pick the one that meets the criteria above.

OpenStudy (anonymous):

D? I am really confused

OpenStudy (radar):

Well lets look at C y-3 = -6(x-2) y-3=-6x +12 y=-6x + 15 Now substitute the value of -2 for x and determine if y=-3

OpenStudy (radar):

Do the same for D. You say you don't want the answer so you need to do the work.

OpenStudy (radar):

Why did you pick D?

OpenStudy (radar):

was it a wag?

OpenStudy (anonymous):

because it has -2 in it, *what is a wag?

OpenStudy (radar):

"Wild A** Guess

OpenStudy (anonymous):

y − 2 = 6(x − 3) y-2=6x-6*3 y-2+6x-18

OpenStudy (anonymous):

D^

OpenStudy (radar):

y - 2 = 6x -18 y = 6x -16 That is D continued ............Is the coefficient of the x a -6 Noooooooo Try another.

OpenStudy (anonymous):

A? y + 3 = −6(x + 2) y+3=-6x+6*2 y+3=-6x+12 y=-6x+9

OpenStudy (radar):

There was an error there, look at the sign for the 12........

OpenStudy (radar):

y + 3 = -6(x + 2) y + 3 = -6X -12 y = -6x - 15 Substitute -2 for the x and solve for y, Y must be equal to -3

OpenStudy (radar):

Did you WAG the A..........you're in luck.

OpenStudy (radar):

Good luck with your studies (and good work)

OpenStudy (anonymous):

so its A, thanks! :)